A conductivity cell having a cell constant of 0.5 cm-1 filled with 0.02 M solution of KCl at 298 K gave a resistance of 20.2 Ω. The water used for preparing the solution had a conductivity of 7.1 × 10–6 S cm–1. Calculate the molar conductivity of 0.02 M KCl solution.


1
Expert's answer
2021-03-25T03:41:29-0400

"\lambda = \frac{\kappa \times 1000}{N}"

"\kappa = \frac{l}{R \times a}"

"\kappa = \frac{0.5}{20.2} = 0.025 \ ohm^{-1} \ cm^{-1}"


"\lambda = \frac{0.025 \times 1000}{0.02} = 1250 \ ohm^{-1} \ cm^2 \ (g-eq)^{-1}"

"=\lambda = 1250 \ ohm^{-1} \ cm^2 \ (g-eq)^{-1}"